How do I find the angles of 5x+3 and 12x+7? They are supplementary angles.
Answer by juaneleanor(3)

Since supplementary angles form 180 degrees

Therefore the equation is…

180=5x+3+12x+7

180=17x+10

Then transpose it …

180-10=17x

170/17=17x/17

10=x

Now that you got the value of x, substitute it…

Angle 1 is 5(10)+3=53 degrees
Angle 2 is 12(10)+7=127 degrees


Eleanor